The results of a survey for an airline are shown below
Traveler Male Female Total
Business 47 72 119
Vacation 71 64 135
Total 118 136 254
Use the chart to find the probability that the traveler was
a) male
b) on vacation given the traveler was female
c) male given the traveler was on vacation

Found 2 solutions by stanbon, nerdybill:
Answer by stanbon(48535)

The results of a survey for an airline are shown below
Traveler Male Female Total
Business 47 72 119
Vacation 71 64 135
Total 118 136 254
Use the chart to find the probability that the traveler was
a) P(male) = 118/254
------------------------------
b) P(on vacation given the traveler was female) = 64/136
------------------------------
c) P(male given the traveler was on vacation) = 71/135
===========================================================
Cheers,
Stan H.

Answer by nerdybill(5404)

The results of a survey for an airline are shown below
Traveler Male Female Total
Business 47 72 119
Vacation 71 64 135
Total 118 136 254
Use the chart to find the probability that the traveler was
a) male
"males on business or vacation"/"total travelers"
(47+71)/(119+135)
118/254 * 100
.465 * 100
46.5%
.
b) on vacation given the traveler was female
"female traveling on vacation"/"total female travelers"
64/(72+64) * 100
= 64/136 * 100
= 47.1%
.
c) male given the traveler was on vacation
"males traveling on vacation"/"total travelers on vacation"
71/(71+64) * 100
= 71/135 * 100
= 52.6%